Shift report 20000714 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Jul 14 2000 - 00:52:51 EDT

  • Next message: http account: "Shift report 20000714 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Jul 14 00:52:51 2000, Shift: 16:00-24:00
    Supervisor: sjs, with: hi
    Log book#2 pages 150 - 151
    
    Shift Summary:
    ------------------------------------------------------------
    Beam in blue ring (at 21:00) followed by beam in
    yellow ring (at 23:40) at injection energy.  Started
    data run 1761 (at injection energy) to check detectors.
    
    KO has moved the operator account from pii3 to opus.
    Going through shift checklist, the change seems to be
    transparent to user..Several changes made to Shifter's
    Checklist (added checks for Si detectors).
    
    C1 and Rich HV off.  Many of the TOFW pedistals seem wide. 
    (see pedistal run 1759).  
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Jul 14 2000 - 00:56:50 EDT